OpenStudy (anonymous):

Given that 3^x=9^y-1, show that x=2y-2

OpenStudy (anonymous):

Its 9^(y-1)

OpenStudy (anonymous):

@tyteen4a03 u r wrong that type of simplification is possible only under product

OpenStudy (anonymous):

3^x=9^(y-1)

OpenStudy (tyteen4a03):

@bhaskarbabu Oh right... my mistake. Represent 9^(y-1) with the base of 3, it becomes 3^(2(y-1)) which in turn becomes 3^(2y-2). The equation becomes 3^x = 3^(2y-2). Now we can reduce the equation to x = 2y-2, which is what you want. Please make sure you use brackets to things people often misunderstand, for example 9^y-1, which is usually interpreted as 9^y minus 1, not 9^(y-1). Thanks :)

OpenStudy (anonymous):

Brilliant thanks.

OpenStudy (anonymous):

@tyteen4a03 oh i too mistaken the question

OpenStudy (anonymous):

How would i reduce.

OpenStudy (anonymous):

@tyteen4a03

OpenStudy (anonymous):

You got your answer..